Please answer all of these questions with clear step and explanations thanks
media%2Fbaf%2Fbaf8f2b2-9991-4b90-b408-11
media%2Ffc8%2Ffc815fa2-6ae8-46a0-87b3-ef
Answer #1

answer1

we need to calculate some financial ratios to explain which firm manages working capital better

WORKING CAPITAL RATIO or CURRENT RATIO = current assets/current liabilities= 3980/3140= 1.27 for coste and4441/4446= 0.99 times for Dila.

This ratio is a measure of the firm's ability to meet its current liabilities.current asets get converted into cash during the operating cycle of the firm and provide the fundsneeded to pay current liabilities.So the higher the current ratio the greater the short term solvency.The current ratio of Costa ltd is higher than Dila.

QUICK RATIO= (current assets less inventories)/current liabilities= 2581/3140= 0.82 times for Coste and 2852/4446= 0.64 for Dila

quick ratio is a fairly stringent measure of liquidity.It is based on those current assets which are highly liquid. Inventories are excluded from the numerator of this ratio, because inventories are deemed to be the least liquid component of current assets.Here also Costa's ratio is better than Dila

COLLECTION RATIO= 365 days * average receiva\bles/revenues

365 * 747/21400= 12.74 days for Coste and 365 * 728/32800 = 8.10 days for DIla.collection period of 8 days rather than 12 days appears to show a prompt collection of receivables and therefore efficient credit management and better working capital management. However this could be because of excessive conservatism in credit granting that may result in loss of some desirable sales. So Coste manages receivables better than Dila

answer 2

allowances/receivables=156/780 * 100 = 20% for 2014 whereas it si 83/830 *100= 10 % for 2015. So we need to take allowances as 20% of receivables in 2015 also .This would be 830 * 20 % =166

The average receivables should appear as 830-166= 664 whereas currently it appears as 747. So additional provision to be made is 747-664= 83

trade receiva\bles figure in balance sheet will be changed to 664

the additional provision of 83 should be taken in the income statement. This will reduce profit to 128-83= 45

It will reduce shareholders equity which will be 6407-83=6324

ans 3

This change will affect collection ratio which will become 365 * 664/21400 =11.32 days

current ratio will be ( 3980-83)/3140 =1.24 times

quick ratio will be( 2581-83)/3140= 0.80

This will not affect answer to part a since the ratios are still favourable than DILa